OpenStudy (madmerc):

Which of the following is the conjugate of a complex number with –1 as the real part and 3i as the imaginary part? -1 - 3i -1 + 3i -3i + 1 3i + 1

OpenStudy (madmerc):

@CShrix

OpenStudy (madmerc):

@Maddy1251 @IHelpYouLearn @Halmos

OpenStudy (madmerc):

@Nnesha @baru @freckles @AlexandervonHumboldt2 @katragaddasaichandra

OpenStudy (lanhikari22):

You know how to write real numbers, right? 5, 2, 12.23323. You know how to write complex numbers too, right? A complex number must contain an imaginary part in it. or in other words, a negative square root! sqrt(-4) is the same as sqrt(-1 * 4) so we just right it as i*sqrt(4) because i = sqrt (-1) and square roots can be broken like: sqrt(2) * sqrt(2) = sqrt(4) = 2. so sqrt(-4) = sqrt(-1) * sqrt(4) = 2i. So how do we put both together, now?

OpenStudy (madmerc):

Put both of what together?

OpenStudy (lanhikari22):

Real, and imaginary numbers. If you want to put both in the same equation, what do you do?

OpenStudy (madmerc):

I don't know...

OpenStudy (madmerc):

Add them?

OpenStudy (lanhikari22):

Yeah, pretty much.

OpenStudy (lanhikari22):

They're not special in that sense, you just add them.

OpenStudy (lanhikari22):

but you keep in mind that, it's like adding this 5 + 6X. so it's absolutely not 11X. it's just... 5 + 6X

OpenStudy (madmerc):

So what am I adding exactly?

OpenStudy (lanhikari22):

Or 6(1 + X) if you've got things like 6 + 6X, but you can't break your ordinary algebra rules. Your question here is asking which of the choices has a real number -1, and a complex number 3i

OpenStudy (lanhikari22):

So the result would be an option that add both.

OpenStudy (madmerc):

so the answer is just -1 + 3i ?

OpenStudy (lanhikari22):

yup

OpenStudy (madmerc):

thank you! can you help me with another?

OpenStudy (lanhikari22):

if it's the same thing, it should follow the same methods, though.

OpenStudy (lanhikari22):

Make sure you understand why, not just how.

OpenStudy (madmerc):

It's not it's \[\frac{ \sqrt{-49} }{ (7 - 2i) - (4 + 9i) }\]

OpenStudy (madmerc):

I don't remember how to do this and I have to finish a bunch of assignments I missed while I was sick

OpenStudy (lanhikari22):

Alright, I'm going to give you the rules. As well as how negative products work. Consider this table, which you might've seen already: i^1 = sqrt(-1) i^2 = -1 i^3 = -i i^4 = 1 Normal intuition would say i^2 = 1, because sqrt(-1) * sqrt(-1) = sqrt(--1) = sqrt(1), right? But, the one thing you need to keep in mind is that INSIDE RADICALS, NEGATIVE/POSITIVE MULTIPLICATION IS NOT -- = + it's because of this:  Basically, inside radicals, negative sign multiplication works 90 degrees instead of 180, which is - to + to - to + sqrt(-) = comes out as i sqrt(--) = sqrt(-^2) you can think of it, and it comes out as... yeah -. or -1 sqrt(---) = sqrt(-^3) = sqrt(-^2 * -) = -sqrt(-) = -i sqrt(----) = sqrt(-^4) = -^2 = -1 * -1 = +1 This should allow you to multiply complex numbers as you want, everything now is according to the normal laws of algebra you use.

OpenStudy (lanhikari22):

Now, for your problem, you can consider i = X, and simplify, while keeping the sign multiplication in mind if needed. In your problem, it's not needed. sqrt(-49) simplifies to norma 7i, so. Good luck!